# Improving User Mental Models of XAI Systems with an Inclusive Design Approach

### v2

@@ Abstract.
Explainable Artificial Intelligence (XAI) systems aim to improve users’ understanding of AI but rarely consider the inclusivity aspects of XAI. Without inclusive approaches, improving explanations might not work well for everyone. This study investigates leveraging users’ diverse problem-solving styles as an inclusive strategy to fix an XAI prototype, with the ultimate goal of improving users’ mental models of AI. We ran a between-subject study with 69 participants. Our results show that the inclusivity fixes increased participants’ engagement with explanations and produced significantly improved mental models. Analyzing differences in mental model scores further highlighted specific inclusivity fixes that contributed to the significant improvement in the mental model. To our surprise, the inclusivity fixes did not improve participants’ prediction performance. However, the fixes did improve inclusivity support for women and promoted equity by reducing the gender gap.
Diversity, Inclusive AI, Explainable AI, Mental Model
copyright: acmcopyrightjournalyear: $9$doi: 99.9999/9999999.9999999journal: TIISpublicationmonth: 99ccs: Human-centered computing Empirical studies in HCI
@@ 1. Introduction
“The curb-cut effect underscores the foundational belief that we are one nation, that we rise or fall together. Without equity, there can be neither progress nor prosperity.” (Blackwell, 2016) —Angela Glover Blackwell, author of the book “Curb-Cut Effect”.
Angela Glover Blackwell’s thoughts on equity resonate with our focus on inclusive XAI. While Blackwell talks about equity, we’re focusing on inclusivity in this paper. As researchers are working to make XAI better, inclusivity asks the question “better for whom?”
By “better”, we mean promoting better mental models of how a user’s AI system is reasoning. In XAI, mental model is defined as the user’s understanding of how an AI system works (Anderson et al., 2020). More formally, mental models are “internal representations that people build based on their experiences in the real world” (Norman and Gentner, 1983). Helping users form suitable mental models is the aim of XAI systems. Although such XAI systems are gaining attention as AI becomes more widespread, the challenges of understanding these systems are also on the rise (Weld and Bansal, 2019).
Improving explanations for one specific user group may not always benefit others if we don’t adopt inclusive approaches in existing AI development practices. One particular inclusive approach involves recognizing users’ diverse problem-solving styles. While demographic diversity is important, considering problem-solving diversity generates actionable design directions (Hamid et al., 2024). Several studies showed how addressing users’ diverse problem-solving approaches reduced inclusivity biases from non-AI applications and systems (Vorvoreanu et al., 2019; Guizani et al., 2022; Burnett et al., 2016). This paper investigates whether focusing on users’ diverse problem-solving styles can reduce inclusivity issues across an information-rich XAI system, ultimately aiming to improve users’ mental models.
To do so, we evaluated how well a group of end users fared in two versions of an XAI prototype. These versions were the work of a team of AI practitioners to address inclusivity issues within their XAI prototype. Their prototype showed a modified version of Tic-Tac-Toe games played sequentially by two AI agents. The AI practitioners leveraged the GenderMag inclusivity method (Hamid et al., 2024; Burnett et al., 2016; Guizani et al., 2022; Vorvoreanu et al., 2019) to find and fix the inclusivity issues. We ran a between-subject study where a group of 34 participants used the “Original” version and 35 participants used the “Post-GenderMag” version. We studied participants’ predictions, their ability to detect flaws in the AI agents, their perception of the AI agents’ behavior, and lastly their mental models. Our investigation centered around these three research questions:
  • RQ1: Did mental models improve with the PostGenderMag version?
  • RQ2: Did prediction performance improve with the PostGenderMag version?
  • RQ3: Did the Post-GenderMag version provide better inclusivity support?
2. Related Works and Background
In recent years, Human-AI Interaction (HAI) research has focused on user performance, system reliability, trustworthiness, and safety (Shneiderman, 2020), while also exploring the dynamics of human and AI control levels (Shneiderman, 2022). The growing body of HAI works aims to increase transparency (Rader et al., 2018), explain AI behaviors (Herlocker et al., 2000; Kulesza et al., 2015), provide guidelines for AI practitioners (Amershi et al., 2019), and address ethical principles in governance (Shneiderman, 2021). Additionally, HAI research advocates for human-centric concerns such as well-being, privacy, and cognitive capacities in the adoption of AI technologies (Ozmen Garibay et al., 2023; Shneiderman, 2022). In our study, we are specifically interested in bringing diverse users’ problem-solving styles together with HAI to improve XAI’s inclusivity.
2.1. Problem-Solving Style and HAI/XAI
There’s been limited investigation into how users’ problem-solving styles influence HAI/XAI. Although AI researchers have examined users’ explanation preferences, they’ve largely overlooked individual differences within these preferences (Dodge et al., 2019). Some research used users’ cognitive diversity to understand why users behave or judge differently when interacting with AI systems. For example, when looking at how fair end-users’ perception of AI was, Dodge et al. (Dodge et al., 2019) assessed the impacts of two factors: i) four types of explanations and ii) data processing. Their findings revealed that participants’ need for cognition significantly impacted their perception of the AI’s fairness. Similarly, Millecamp et al.’s work (Millecamp et al., 2020) found that diverse end-users’ cognitive styles showed differences in their trends to interact with the AI’s explanations. This investigation discussed that the variation in the definitions of cognitive styles did not have much effect and any of the definition variations could help identify valuable insights.
Goldberg et al. (Goldberg, 1990) presented an alternative approach to capturing individual differences through five personality traits— Openness, Conscientiousness, Extraversion, Agreeableness, and Neuroticism (OCEAN) (Goldberg, 1990). They found that one of the personality traits, “Openness” influenced users’ interaction with the interface. Jiang et al. (Jiang et al., 2000) also used a personality trait, self-confidence to evaluate participants’ acceptance of the expert system’s output. Other researchers also used some form of problem-solving styles, cognitive styles, and personality traits to explain user behavior (Giboney et al., 2012; Kulesza et al., 2012; Felmingham et al., 2021; Djulbegovic et al., 2014). However, none of these works considered a wider range of problem-solving styles as we did in our study.
In this study, we used problem-solving styles defined by users’— attitude towards risk, computer self-efficacy, motivations, information processing style, and learning style.
2.2. GenderMag
@@
We used the GenderMag method to define the end users’ five problem-solving style ranges used in this study. GenderMag, also known as the Gender-Inclusiveness Magnifier, is a usability inspection method to find and fix gender-inclusivity bugs in problem-solving software (Hamid et al., 2024; Burnett et al., 2016; Guizani et al., 2022; Vorvoreanu et al., 2019). This method focuses on five facets of gender differences (statistical), brings them to life through three faceted personas, and encapsulates the use of these facets into a systematic process through a gender specialization of the Cognitive Walkthrough (CW) (Wharton et al., 1994). Software practitioners have used this method to evaluate the software they design and develop from a gender-inclusiveness perspective (Burnett et al., 2016; Guizani et al., 2022; Vorvoreanu et al., 2019).
@@
The GenderMag method’s five facets capture five factors in an individual’s problem-solving style: their motivations for using technology, information processing styles, computer self-efficacy, attitude toward risk in using unfamiliar technology features, and learning style (by process vs. by tinkering) (Burnett et al., 2016; Hamid et al., 2024). The “inclusivity bugs” uncovered by GenderMag are instances where a technology product fails to fully accommodate the complete range of values associated with the five facets, leading to a disproportionate impact on individuals whose problem-solving styles are unsupported. These bugs are also considered gender-inclusivity issues because the facets reflect statistical differences among gender preferences in problem-solving styles (Burnett et al., 2016; Guizani et al., 2022; Vorvoreanu et al., 2019; Hamid et al., 2024). GenderMag reveals inclusivity issues that are statistically associated with gender preferences. GenderMag does not use peoples’ gender identities to find inclusivity concerns, it uses only their problem-solving styles.
@@
The GenderMag method has three personas: Abi (Abigail/Abishek), Pat (Patricia/Patrick), and Tim (Timara/Timothy) (see Table 1). Statistically, the problem-solving styles of Abi disproportionately skew toward women, for Tim they disproportionately skew toward men, and Pat provides a third set of values within Abi and Tim (Burnett et al., 2016; Hamid et al., 2024).
@@
Abigail/Abishek (”Abi”) Patricia/Patrick (”Pat”) Timara/Timothy (”Tim”)
@@
The GenderMag faceted personas are integrated into a systematic specialized Cognitive Walkthrough process. A group of evaluators adopt a specific GenderMag persona, say Abi, customize the persona’s background (optional), and systematically “walkthrough” each step of a use case. At each step, they answer three questions: i) will ¡Abi¿ have formed this subgoal as a step to their overall goal? why?, ii) will ¡Abi¿ take the action the developers expected at this step? Why?, and iii) if ¡Abi¿ took the action, will they know that they did the right thing and are making progress toward their goal? Why? If the evaluators find any answer is negative because of Abi’s particular problem-solving styles, they report an inclusivity bug, provide their reasons, and note down the problem-solving styles that support their reasonings. Later, when fixing an inclusivity bug, these notes ensure that the fixes are designed around Abi’s problem-solving styles that highlighted the bug.
The evaluators relevant to our study were a group of AI practitioners whom we called “Team Game” in this paper. Team Game evaluated their own XAI prototype from the perspective of the Abi persona. They chose the Abi persona as Abi represents the group of users who are most often left out of the development process which results in more barriers for these users. Team Game’s before GenderMag and after GenderMag version of the prototype were then used in our study.
2.3. XAI in Sequential Domains
The AI practitioners of our study designed their XAI prototype specifically for a sequential domain. A sequential domain refers to a specific situation or area where actions or events unfold in a particular order, following a series of steps or stages in a predetermined sequence. For instance, consider a board game like Chess, where players take turns making moves and the progression of the game follows an ordered sequence of moves. In recent years, there has been a growing interest in developing AI systems that can play sequential games like Chess, Shogi, and Go. This has led to the development of AI players that can compete with and even outperform the world champions of these games (Silver et al., 2016, 2018). Real-Time Strategy (RTS) games have also seen similar achievements (Vinyals et al., 2019b, a).
Despite such progress, explaining AI players’ actions in sequential domains remains a major challenge. In an attempt to understand this challenge, Metoyer et al. (Metoyer et al., 2010) studied how experienced users explain strategies and gameplay in RTS games. In another study, Penney et al. (Penney et al., 2018) found that even experienced users encounter challenges when seeking information about an intelligent agent in an RTS setting. One result was that participants showed a preference for “What” information over “Why”. In a sequential game domain, understanding the causal relationships among relevant variables within the game can also help users (Madumal et al., 2020). As Madumal et al. (Madumal et al., 2020) showed offering causal explanations of AI agents can enhance participants’ ability to predict tasks, as well as increase their satisfaction and trust. In another study, Dodge et al. (Dodge et al., 2022) observed that when ranking AI agents playing a sequential game like Tic-Tac-Toe, participants highlighted the importance of detailed explanations to differentiate between closely matched agents. That study also noted that participants showed a preference for receiving multiple explanations over a single one when making decisions.
2.4. Mental Models and Explainable AI (XAI)
The opaque nature of AI systems poses a challenge for end-users with low or no expertise in AI (Wortham et al., 2017; Tsai and Carroll, 2022). This especially hinders users’ ability to seamlessly interpret and trust the system’s output (Kulesza et al., 2015). Lim et. al (Lim et al., 2009) regarded such lack of clarity as an intelligibility concern for AI systems.
To improve the clarity of an AI system, researchers and developers often provide explanations of the system to its end-users (Alipour et al., 2021; Brachman et al., 2023; Schraagen et al., 2020; Pierson et al., 2024). According to Hoffman et al., explanations that are effective and satisfactory for users can facilitate the development of a good mental model (Hoffman et al., 2019). Kulesza et al. demonstrated that providing explanations about the debugging process of a machine-learning system improved participants’ understanding of the system (Kulesza et al., 2015, 2010). Also, the explanations improved their participants’ ability to rectify errors. Other studies also found additional benefits of adding explanations such as improvements in users’ self-reported satisfaction (Kapoor et al., 2010; Amershi et al., 2014) and feelings of trust towards the AI systems (Lim et al., 2009; Tsai et al., 2021; Schraagen et al., 2020).
Mental models have been captured using different methods including questionnaires about the agent, mental model drawings, and task performance (Brachman et al., 2023; Schraagen et al., 2020; Zhang et al., 2024; Bansal et al., 2019; Alipour et al., 2021). Hoffman et al. (Hoffman et al., 2019) proposed measures that evaluate the quality, user satisfaction, and user understanding of the explanations. According to Hoffman et al., mental models can be deduced from empirical evidence, even though individuals might find it challenging to fully articulate their understanding (Hoffman et al., 2019). However, the authors suggested that through guided task reflection, people can effectively communicate their understanding of an event or system, including its fundamental concepts and principles. The authors provided a list of 11 methods that researchers have used to derive mental models across various domains, including cognitive psychology, expert system knowledge acquisition, instructional design, and educational computing. From these 11 methods, Anderson et al. used the “Prediction Task” and “Task Reflection” to understand users’ mental models of a Reinforcement Learning agent (Anderson et al., 2019). In our paper, we have utilized these two methods and also integrated the “Glitch Detector Task”, and “Self-Explanation Task” from Hoffman et al.’s 11 methods to know if users were able to find any flaws in the given explanations and their own understanding of the AI agent.
Thus, although HAI and XAI research have made significant progress in exploring users’ mental models, they often overlooked the potential impact of users’ diverse problem-solving styles on their overall understanding of AI systems. Recognizing this gap, we wanted to check the effectiveness of GenderMag-driven cognitive walkthroughs that put users’ diverse problem-solving styles at the center stage of finding and fixing inclusivity issues. Historically, HCI practitioners leveraged GenderMag for a wide variety of user experience improvements such as inclusivity bug removal (Guizani et al., 2022), task performance improvement (Vorvoreanu et al., 2019), customer rating improvement (Burnett et al., 2017), etc. However, no prior GenderMag research has evaluated GenderMag’s applicability and effectiveness within AI-powered systems. Addressing this gap in GenderMag studies, our goal is to investigate how successful GenderMag fixes can be in improving users’ engagement and their mental models of AI.

### v3

# Inclusive Design of AI’s Explanations: Just for Those Previously Left Out, or for Everyone?

@@ Abstract.
Motivations: Explainable Artificial Intelligence (XAI) systems aim to improve users’ understanding of AI, but XAI research shows many cases of different explanations serving some users well and being unhelpful or worse to other users. With non-AI systems, some software practitioners have addressed problems like this by using inclusive design approaches. In some fortuitous cases, their improvements turned out to be “curb-cut” improvements—not only addressing the needs of underserved users, but also making the products better for everyone else as well. This raises the possibility that if AI practitioners used inclusive design approaches to improve their AI product’s explanations, they too might create curb-cut improvements, i.e., better explanations for everyone.
Objectives: To consider this possibility, we investigated the effects of inclusivity-driven fixes in an XAI prototype. The prototype and fixes came from an AI product team who had adopted an inclusive design approach (GenderMag) to improve their XAI prototype. Our objective was to investigate the curb-cut effects of the AI team’s inclusivity-driven fixes on users’ mental models of AI.
Methods: We ran a between-subject study with 69 participants who had no AI background. 34 participants used the original version of the XAI prototype and 35 used the version with the AI team’s inclusivity fixes. To understand the curb-cut effects of inclusivity-driven fixes in the XAI prototype, we compared the two groups’ mental model concepts scores, prediction accuracy, and inclusivity.
Results: Our investigation produced four main results. First, it revealed several curb-cut effects of the AI team’s inclusivity fixes, such as overall increased engagement with explanations and better mental model concepts scores. These improvements revealed specific inclusivity fixes with curb-cut properties. However (second), the inclusivity fixes did not improve participants’ prediction accuracy scores—instead, it appears to have harmed them. This “curb-fence” effect (opposite of the curb-cut effect) on participants’ prediction accuracy scores revealed the AI explanations’ double-edged impact, sometimes operating for better and sometimes unexpectedly for worse. Third, the results showed the AI team’s inclusivity fixes brought significant improvements for users whose problem-solving styles had previously been underserved. Further (fourth), the AI team’s fixes reduced the gender gap by 45%.
Problem-Solving Style, Inclusive Design, Explainable AI, Mental Model
copyright: acmcopyrightjournalyear: $9$doi: 99.9999/9999999.9999999journal: TIISpublicationmonth: 99ccs: Human-centered computing Empirical studies in HCIccs: Human-centered computing User studiesccs: Computing methodologies Intelligent agents
@@ 1. Introduction
Figure 1. A sidewalk’s curb cut.
Consider curb cuts in sidewalks, like the one in Figure 1. These cut-out sections of sidewalks were designed for a population who was not well-served by sidewalks before curb cuts, namely people in wheelchairs. And they did make sidewalks better for people in wheelchairs—but also, for everybody else (e.g., parents with strollers, people pushing heavy carts, travelers with wheeling suitcases, etc.) (Blackwell, 2016). This property of curb cuts has been termed the curb-cut effect: an effect of a design choice made to better serve an underserved population that also turns out to help everyone else too, even those who were already well-served (Blackwell, 2016).
In the field of Explainable AI (XAI), many researchers have reported problems with explanations’ effectiveness for diverse users in diverse situations (Anderson et al., 2020; Bhatt et al., 2020; Gajos and Mamykina, 2022; Gerlings et al., 2022; Rosenfeld and Richardson, 2019; Wang et al., 2024). This situation not only suggests (1) the need to attend to users’ diversity, but also (2) ongoing problems generally with the effectiveness of AI explanations. In this paper, we consider these two problems as one, to investigate the possibilities of curb-cut effects. Specifically, can an inclusive design approach produce explanations that not only help some underserved group(s) of users, but also are “just plain better” explanations for everyone?
For our paper’s XAI scope, by “better”, we mean promoting more accurate mental models of how a user’s AI system is reasoning. In XAI, mental models are defined as the user’s understanding of how an AI system works (Anderson et al., 2020). Helping users form suitable mental models is an objective of many XAI systems (Kulesza et al., 2013; Nourani et al., 2021; Schraagen et al., 2020).
Thus, this paper investigates if inclusive design can bring curb-cut effects to XAI’s explanations in the form of better mental models for everyone. To do so, we evaluated how well two groups of end users fared in two versions of an XAI prototype. These versions were the work of a team of AI practitioners (Team Game) who were hoping to address inclusivity issues within their XAI prototype. Their prototype was a modified version of Tic-Tac-Toe games played sequentially by a red AI agent and a blue AI agent. The prototype was supported by several kinds of explanations of the blue AI agent’s behaviors. Prior to our study, the AI practitioners had used the GenderMag inclusivity method to find and fix inclusivity issues in their prototype (Anderson, 2024).
GenderMag  (Burnett et al., 2016) is an inclusivity evaluation method that considers five ranges of users’ problem-solving style types: the ranges of users’ various Motivations for using technology, of their diverse Information Processing Styles, of their levels of Computer Self-Efficacy, of their Attitudes Toward Risk in using unfamiliar technology features, and of their Learning Styles (Hamid et al., 2024). Several studies have shown GenderMag’s effectiveness in non-AI systems (e.g., (Burnett et al., 2016; Guizani et al., 2022; Burnett et al., 2017)), as we explain further in Section 2.1.
For the current paper, we ran a between-subject study where a group of 34 participants used the “Original” version and 35 participants used the “Post-GenderMag” version. Since the prototype is attempting to explain to end users, our participants were restricted to only those with no AI background. For the “better for everyone” aspect of curb-cut effects, we measured the Original participants’ vs. the Post-GenderMag participants’ conceptual understanding of the AI agents’ behavior, their ability to predict the blue AI agent’s next move, and their engagement with the system’s explanations. For the “better for the underserved” aspect of curb-cuts, we also disaggregated these measures by users’ problem-solving style values and by their genders.
Our investigation centered around these three research questions:
  • RQ1: Did the PostGenderMag participants have better mental model concepts scores than the Original participants?
  • RQ2: Did the PostGenderMag participants have better prediction accuracy than the Original participants?
  • RQ3: Was the Post-GenderMag prototype more inclusive than the Original prototype?
2. Background and Related Works
2.1. Background: GenderMag
@@
The AI practitioners team used the GenderMag inclusivity method, so our results rest on GenderMag concepts. GenderMag (Gender-Inclusiveness Magnifier) is an inspection method to find and/or fix inclusivity bugs in problem-solving software (Hamid et al., 2024; Burnett et al., 2016; Guizani et al., 2022; Vorvoreanu et al., 2019). The method focuses on diversity of problem-solving styles (“facets” in GenderMag terminology) which statistically cluster by gender. The method uses these facets as its core, brings them to life through three faceted personas, and encapsulates the use of these facets into a systematic process through a specialized Cognitive Walkthrough (CW) (Wharton et al., 1994). Numerous software practitioners have used GenderMag to evaluate and/or fix the inclusiveness of their software products (Burnett et al., 2017; Cunningham et al., 2016; Guizani et al., 2022; Hilderbrand et al., 2020; Kirillova and Malikh, 2017; Murphy-Hill et al., 2024; Santos et al., 2023; Shekhar and Marsden, 2018; Vorvoreanu et al., 2019).
@@
The GenderMag method’s five facets capture five factors in an individual’s problem-solving approach: their Motivations for using technology, Information Processing Styles, Computer Self-Efficacy, Attitude Toward Risk in using unfamiliar technology features, and Learning Style (by Process vs. by Tinkering) (Burnett et al., 2016). The “inclusivity bugs” uncovered by GenderMag are instances where a technology product fails to fully accommodate the complete range of values associated with the five facets, leading to a disproportionate impact on individuals whose problem-solving styles are unsupported. These bugs are also considered gender-inclusivity issues because the facets reflect statistical differences among gender preferences in problem-solving styles (Burnett et al., 2016; Guizani et al., 2022; Vorvoreanu et al., 2019; Hamid et al., 2024). GenderMag does not use peoples’ gender identities to find inclusivity issues, it uses only their problem-solving styles.
@@
The GenderMag method represents the range of facet values with three personas: Abi (Abigail/Abishek), Pat (Patricia/Patrick), and Tim (Timara/Timothy) (see Table 1). Statistically, the problem-solving styles of Abi disproportionately skew toward women, for Tim they disproportionately skew toward men, and Pat provides a third set of values (Burnett et al., 2016; Hamid et al., 2024).
@@
Abigail/Abishek (“Abi”) Patricia/Patrick (“Pat”) Timara/Timothy (“Tim”)
@@
The GenderMag faceted personas are integrated into a systematic specialized Cognitive Walkthrough process. A group of evaluators adopt a specific GenderMag persona, say Abi, customize the persona’s background (optional), and systematically “walkthrough” each step of a use case. At each step, they answer three questions: (i) will ¡Abi¿ have formed this subgoal as a step to their overall goal? why?, (ii) Will ¡Abi¿ know what to do at this step? Why?, and (iii) if ¡Abi¿ does the right thing, will they know that they did the right thing and are making progress toward their goal? Why? If the evaluators find any answer is negative because of the persona’s particular problem-solving styles, they report an inclusivity bug, provide their reasons, and note down the problem-solving styles that support their reasoning. When fixing the bugs, these notes help ensure that the fixes are designed around the persona’s problem-solving styles that highlighted the bugs.
2.2. Related Work
XAI researchers have investigated different explanation properties’ abilities to improve users’ mental models and understanding of AI systems. One group of XAI researchers examined the effect of parsimony and stochasticity of an AI’s error boundary on users’ mental models of AI capabilities (Bansal et al., 2019). They found that simpler and more predictable error boundaries improved users’ collaboration with AI and AI-assisted decision-making tasks (Bansal et al., 2019). Tsai et al. showed another XAI strategy—interweaving an explanation in the conversation flow of an intelligent online symptom checker—that improved users’ understanding of the diagnoses and medical recommendations (Tsai et al., 2021). A common XAI research area involves comparing different types of explanations to understand each explanation’s impacts on users’ mental models of AI (Alipour et al., 2021; Tsai et al., 2021; Wang and Yin, 2021). For example, Wang and Yin (Wang and Yin, 2021) studied the effects of four model-agnostic explanations (feature importance, feature contribution, nearest neighbors, and counterfactual) on participants’ understanding of the Machine Learning (ML) model. They found two effects in tasks where participants perceived themselves with higher domain expertise: (i) feature importance and counterfactual explanations increased participants’ objective understanding; and (ii) all four model-agnostic explanations increased their subjective understanding (Wang and Yin, 2021). In contrast, when participants were less knowledgeable about the task, explanations did not improve their understanding, uncertainty awareness, or trust. In a similar study, Cai et al. found that normative explanations (i.e., showing what is “normal”) improved user understanding and trust in an ML system, especially when the algorithm failed. In contrast, although comparative explanations revealed algorithmic limitations, they did not always improve users’ understanding of the algorithm (Cai et al., 2019).
When studying the effects of multiple explanations, XAI researchers also suggested providing more than one explanation to improve users’ mental models. Anderson et al. found that participants who used multiple explanations (saliency maps and reward-decomposition bars) had better mental model scores compared to participants who saw one or no explanations (Anderson et al., 2020). This result is similar to Schraagen et al.’s study where mixed explanations—combining both reasons and causes—helped participants achieve the best understanding of a self-driving car’s decisions (Schraagen et al., 2020). Dodge et al. also proposed a hybrid approach to explain ML systems to support better human-in-the-loop workflow, where global explanations would help users understand and evaluate the ML system and local explanations would help analyze individual predictions (Dodge et al., 2019).
To improve users’ interactions and understanding of AI, a few XAI researchers have investigated human attributes like cognitive biases, need for cognition (NFC), and user diversity. For example, Nourani et al. (Nourani et al., 2021) investigated the effects of anchoring bias, a form of cognitive bias, on users’ mental model formation in XAI systems. They found that participants who observed system strengths first were more prone to anchor bias and made more errors, but developed a more accurate mental model of the system than the participants who noticed the weaknesses first. In another study, Millecamp et al. examined the effects of users’ NFC on their interactions, trust, and decision-making in a recommender system through the use of explanations (Millecamp et al., 2020). They found that participants with high NFC actively sought and engaged with explanations to control and steer the recommendation process whereas participants with low NFC sought explanations when they encountered unexpected recommendations or when searching for specific content (Millecamp et al., 2020). Similarly, Kulesza et al. studied the mental demands of explanation on users and found that explanations must balance both soundness and completeness to effectively enhance user’s understanding of the AI systems, as overly simplistic or incomplete explanations can lead to cognitive overload and diminished trust (Kulesza et al., 2013). A recent human-centric approach in XAI involves integrating social explanations that use socio-contextual cues during the explanation process to provide actionable insights to users (Gong et al., 2024; Ehsan et al., 2021; Liao and Varshney, 2021; Sun et al., 2022). For instance, Brachman et al. found that social explanations, which offered insights into other users’ actions, were more helpful compared to system-focused explanations that provided details about the system’s input processing and output generation (Brachman et al., 2023). However, the participants’ mental models did not differ based on the explanation types.
GenderMag is centrally about inclusivity, and many HCI practitioners have used GenderMag to evaluate and/or improve non-AI products’ inclusivity. Examples include using GenderMag to evaluate or improve inclusivity of Open Source Software project sites (Guizani et al., 2022; Santos et al., 2023), online courseware (Chatterjee et al., 2022), a system for IT professionals (Burnett et al., 2017), learning management platforms (Hilderbrand et al., 2020; Shekhar and Marsden, 2018), education software for middle-schoolers (Burnett et al., 2016), and a code review tool (Murphy-Hill et al., 2024). A few researchers have also begun using GenderMag’s problem-solving styles for measurement within AI contexts. Nam et al. used the problem-solving styles, specifically Information Processing Style and Learning Style, to understand how users interacted with a Large Language Model (LLM) designed to help with code understanding (Nam et al., 2024). They found that participants’ usage of AI features varied with their Learning Style. For example, participants with a process-oriented Learning Style were more likely to use the “Followup” AI feature than the tinkerers. Researchers have also used problem-solving styles to measure AI products’ inclusivity. Anderson et al. compared the inclusiveness of 16 AI products designed using Amershi et al.’s guidelines (Amershi et al., 2019) against versions of those products that violated those guidelines, and found that the guideline-compliant AI products were more inclusive for users with diverse problem-solving styles than those violating the guidelines (Anderson et al., 2024). They also found that supporting diverse problem-solving styles improved the demographic inclusivity of AI products. Finally, Vorvoreanu et al. used GenderMag to evaluate and improve an AI-powered search tool (Vorvoreanu et al., 2019). However, previous research has not investigated GenderMag, or any other inclusive design method, as a vehicle for improving users’ mental models of AI systems.